\begin{thebibliography}{30}

\vspace{-0.3cm}
\bibitem{Lif2019}
%7
V.~I. Burenkov and E. Liflyand, Hausdorff operators on Morrey-type spaces, Kyoto J. Math. {\bf 60} (2020), no.~1, 93--106.

\vspace{-0.3cm}
\bibitem{Bracewell1965}
%5
R.~N. Bracewell, {\it The Fourier transform and its applications}, McGraw-Hill, New York-Toronto-London, 1965.

\vspace{-0.3cm}
\bibitem{Loukas2024}
%6
L. Grafakos, {\it Fundamentals of Fourier Analysis}, Graduate Texts in Mathematics, 302, Springer, Cham, 2024.

\vspace{-0.3cm}
\bibitem{Lu1995}
%1
S.~Z. Lu, {\it Four lectures on real $H^p$ spaces}, World Sci. Publ., River Edge, NJ, 1995.

\vspace{-0.3cm}
\bibitem{Liflyand2021}
%3
A. Debernardi~Pinos and E. Liflyand, Approximation via Hausdorff operators, Canad. Math. Bull. {\bf 64} (2021), no.~3, 512--529.

\vspace{-0.3cm}
\bibitem{Stein1993}
%2
E.~M. Stein, {\it Harmonic analysis: real-variable methods, orthogonality, and oscillatory integrals}, Princeton Mathematical Series Monographs in Harmonic Analysis, 43 III, Princeton Univ. Press, Princeton, NJ, 1993.

%\vspace{-0.3cm}
%\bibitem{Vinogradov2023}
%%4
%O.~L. Vinogradov, Order of approximation by partial Hausdorff integrals, J. Math. Anal. Appl. {\bf 522} (2023), no.~1, Paper No. 126924, 10 pp.






\bigskip\medskip

\noindent Zifei Yu and Baode Li (Corresponding author),
\medskip

\noindent College of Mathematics and System Sciences\\
Xinjiang University\\
Urumqi, 830017\\
P. R. China
\smallskip

\noindent{E-mail }:\\
\texttt{107552300599@stu.xju.edu.cn} (Zifei Yu)\\
\texttt{baodeli@xju.edu.cn} (Baode Li)\\
\bigskip \medskip

\end{thebibliography}

\documentclass[11pt, twoside]{article}
\usepackage{amsfonts}
\usepackage{ulem}

\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{amsthm}
\usepackage{xcolor}
\usepackage{mathrsfs}
\usepackage{mathabx}

\usepackage{verbatim}

\allowdisplaybreaks

\allowdisplaybreaks

\def\R{{\mathbb R}}

\def\R{{\mathbb R}}

\def\z+{{\mathbb Z}_+}

\def\bint{{\ifinner\rlap{\bf\kern.30em--}
\int\else\rlap{\bf\kern.35em--}\int\fi}\ignorespaces}

\def\sbint{{\ifinner\rlap{\bf\kern.32em--}
\hspace{0.078cm}\int\else\rlap{\bf\kern.45em--}\int\fi}\ignorespaces}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}

\newtheorem{proposition}[theorem]{Proposition}
\theoremstyle{definition}
\newtheorem{example}[theorem]{Example}

\newtheorem{definition}[theorem]{Definition}
\numberwithin{equation}{section}

\numberwithin{equation}{section}

\numberwithin{equation}{section}

\begin{document}

\arraycolsep=1pt

\title{\Large\bf Approximation via partial Hausdorff integrals on $H^1(\R)$ \footnotetext{\hspace{-0.35cm} {\it 2020
Mathematics Subject Classification}. {41A25, 42A38.}
\endgraf{\it Key words and phrases.} Hausdorﬀ operators, approximation, Hardy space.
\endgraf This project is supported by Natural Science Foundation of Xinjiang Province (No.2024D01C40) and National Natural Science Foundation of China (No.12261083).
\endgraf $^\ast$\,Corresponding author.
}}
\author{Zifei Yu and Baode Li$^\ast$}
\date{ }
\maketitle

\vspace{-0.8cm}

\begin{center}
\begin{minipage}{13cm}\small
{\noindent{\bf Abstract.}
We obtain the result of approximating \( f \) in the \( H^1(\mathbb{R}) \) norm using partial Hausdorff integrals. Specifically, by leveraging the homogeneous multiplier theory of \( H^1(\mathbb{R}) \) and the \( K \) functional theory, one result from Pinos and Liflyand [CMB,~2021,~64,~no.3] is extended from \( L^p(\mathbb{R}) \) ( \( 1 \leq p \leq \infty \)) to \( H^1(\mathbb{R}) \). As applications, four examples of partial Hausdorff integrals are also given.} 
\end{minipage}
\end{center}

\section{Introduction\label{s1}}

Let $\varphi$ be a measurable real-valued function on $\R$. Assume that $a:\R\to\R$, $a$ is measurable, $a\neq0$ a.e., and for every set $E$ of zero Lebesgue measure, the set $a^{-1}(E)$ also has zero measure. Under these assumptions, Burenkov and Liflyand \cite[Theorem 1]{Lif2019} proved that the function \((x,t) \mapsto f(a(t)x)\) is measurable on $\R^2$. Let $\varphi|a|^\frac{1}{2}\in L^1(\R)$. A general {\it Hausdorff operator} $H$ of $f\in L^2(\R)$ is defined by
\begin{equation}\label{eq23}
H(f)(x):=H_{\varphi,a}(f)(x)=\int_{\R} \varphi(t)|a(t)|f(a(t)x)dt,~x\in \R.
\end{equation}
By Minkowski's inequality and substituting $a(t)x=\tilde{x}$, it is easy to verify that $H$ is bounded on $L^2(\R)$:
\[
\|Hf\|_{L^2}
\leq \int_{\R} |\varphi(t)||a(t)| \left(\int_{\R}|f(a(t)x)|^2 dx\right)^\frac{1}{2}dt=\int_{\R}|\varphi(t)||a(t)|^\frac{1}{2}dt\|f\|_{L^2}.
\]

Taking $a(t)=1/t$ when $t\neq0$ and $a(0)=0$, then $H$ is a {\it one-dimensional Hausdorff operator} $H_\varphi$, i.e.,
\[
H_{\varphi}(f)(x):=\int_{\R} \frac{\varphi(t)}{|t|} f\left(\frac{x}{t}\right) dt.
\]

Suppose $a$ is additionally odd and such that $|a|$ is decreasing, positive and bijective on $(0,+\infty)$. Pinos and Liflyand \cite{Liflyand2021} defined 
\begin{align*}
({H_N\widehat{f}})\widecheck{~}(x)
&:=
\frac{1}{2\pi}\int_{-N}^{N} H\widehat{f}(u)e^{iux}du
\\
&=\frac{1}{2\pi}\int_{-N}^{N} \int_{\R}\varphi(t)|a(t)|\int_{\R}f(s)e^{-ia(t)su}dsdte^{iux}du
\\
&=\frac{1}{\pi}\int_{\R}\varphi(t)|a(t)|\int_{\R}f(s)\frac{\sin N(x-a(t)s)}{x-a(t)s}ds dt, \quad N>0.
\end{align*}
By substituting $\tilde{s}=N(\frac{x}{a(t)}-s)$,
\begin{equation}\label{eq25}
(H_N\widehat{f})\widecheck{~}(x)=\frac{1}{\pi}\int_{\R}\varphi(t)\int_{\R}f\left(\frac{x}{a(t)}-\frac{s}{N}\right)\frac{\sin|a(t)|s}{s}dsdt.
\end{equation}
Suppose that $\varphi\in L^1(\R)$ and $\int_{\R}\varphi(t)dt=1$. To approximate \( f \), Pinos and Liflyand \cite{Liflyand2021} defined the {\it  partial Hausdorff integrals} by
\begin{align}
F_N(x):&=\left(H_N\widehat{f(x+\cdot)}\right)\widecheck{~}(0)=\left(H_N\widehat{\tau_xf}\right)\widecheck{~}(0)\nonumber\\
&=\frac{1}{\pi}\int_{\R}\varphi(t)\int_{\R}f\left(x-\frac{s}{N}\right)\frac{\sin(|a(t)|s)}{s} {d}s dt.\label{eq24}
\end{align}
In \cite{Liflyand2021}, Pinos and Liflyand obtained the approximation of \( f \) by \( F_N \) in the \( L^p(\R) \)-norm using the {\it \( L^p(\R) \)-modulus of continuity} when $1\leq p\leq\infty$, $\varphi\max\{|a|^\frac{1}{p},|a|^\frac{1}{2}\}\in L^1(\R^n)$ and $f\in L^1\cap L^p(\R^n)$.

In our proof, for convenience, we set \( 1/N = \epsilon \) in $(\ref{eq24})$ and denote
\begin{equation}\label{eq11}
F_\epsilon(x)
:=
\frac{1}{\pi}\int_{\R}\varphi(t)\int_{\R}f(x-\epsilon s)\frac{\sin(|a(t)|s)}{s} ds dt. 
\end{equation}
We obtained the approximation of \( f \) by \( F_\epsilon \) in the \( H^1(\R) \)-norm using the {\it K functional} from \cite[Chapter 4]{Lu1995}.

The structure of this paper is as follows: In Section \ref{s2}, we introduce the necessary definitions and notations; in Section \ref{s3}, we first prove that the partial Hausdorff integrals are uniformly bounded on \( H^1(\R) \), and then we show that \( f \) can be approximated by the partial Hausdorff integrals under the \( H^1(\R) \)-norm, along with several necessary lemmas; and in Section \ref{s4}, we provide four examples.

The proof of  \cite[Theorem 2.2]{Lif2019} uses the modulus of continuity of \( L^p(\mathbb{R}^n) \) (\( 1 \leq p \leq \infty \)) to show that partial Hausdorff integrals can approximate \( f \in L^p(\mathbb{R}^n) \) under the \( L^p(\mathbb{R}^n) \)-norm. Different from this approach, we utilize the $K$ functional to prove that, under the same conditions, partial Hausdorff integrals can approximate \( f \in H^1(\mathbb{R}) \) under the \( H^1(\mathbb{R}) \)-norm.




\section{Notations and definitions \label{s2}}

\begin{definition}\label{def2.2}\cite{Loukas2024}
For \(1\leq p<\infty\), we define the \(L^p(\R)\) norm of a measurable function \(f\) by
\[
\|f\|_{L^p(\R)}:=\left(\int_{\R}|f(x)|^p d x\right)^{\frac{1}{p}}.
\]
For any \(1\leq p< \infty\), we define \(L^p(\R)\) to be the space of all measurable functions \(f\) with \(\|f\|_{L^p(\R)}<\infty\). 

\end{definition}

\begin{definition}\cite{Loukas2024}
Let \( f \in L^2(\mathbb{R}) \). The {\it Fourier transform} \( \widehat{f} \) of \( f \) is defined by
\[
\widehat{f}(x)=\int_{\R} f(x) e^{-ixy} dy\overset{L^2(\R)}{:=}\lim_{R\to+\infty}\int_{|x|\leq R} f(x)e^{-ixy}dy.
\]
Analogously define the {\it inverse Fourier transform} \( f^\vee \) of $f$ by
\[
f^\vee(x)=\frac{1}{2\pi}\int_{\R}f(x)e^{ixy}dx\overset{L^2(\R)}{:=}\lim_{R\to+\infty}\frac{1}{2\pi}\int_{|x|\leq R} f(x)e^{ixy}dx.
\]
\end{definition}

\begin{definition}
Suppose \(\Phi\in \mathscr{S}\left(\R\right)\) with $\int_{\R} \Phi(x) d x \neq 0 .$
For any tempered distributions $f\in\mathscr{S}'(\R)$, the {\it radial maximal operator} $M^+_{\Phi}$ of $f$ is defined by
\[
M_{\Phi}^+f(x):=\sup _{0<s<\infty}\left|\Phi_s * f(x)\right|,\quad x\in\R,
\]
where \(\Phi_s(x):=\frac{1}{s} \Phi\left(\frac{x}{s}\right)\).
The {\it Hardy space} \(H^1(\R)\) is the space of all tempered distributions \(f\) satisfying
\[
\|f\|_{H^1}:=\left\|M_{\Phi}^+f\right\|_{L^1(\R)}<\infty .
\]
\end{definition}

\begin{definition}\label{def1}\cite[Page 174]{Lu1995}
Suppose $f\in H^1(\R)$ and $\sigma>0$. The {\it $\sigma$-th order Riesz derivative} $I^\sigma f$ is defined by
\begin{equation*}\label{eq1}
\widehat{I^\sigma f}(x)=|x|^\sigma \widehat{f}(x).
\end{equation*}
Then we define some subspaces $H^{1,\sigma}(\R)$ of $H^1(\R)$ as follows
\begin{equation*}\label{eq2}
H^{1,\sigma}:=\{f\in H^1: I^\sigma f\in H^1\}.
\end{equation*}
\end{definition}

\begin{definition}\label{def2}\cite[Page 174]{Lu1995}
Suppose that $\sigma>0$, $t>0$ and $f\in H^1(\R)$. {\it The $\sigma$-th order $K$ functional of $f$} is defined by
\begin{equation*}
K_\sigma(f,t)_{H^1}:=\inf_{g\in H^{1,\sigma}}\{\|f-g\|_{H^1}+t^\sigma\|I^\sigma g\|_{H^1}\}.
\end{equation*}
\end{definition}

\begin{proposition}\label{pro1}\cite[Page 174]{Lu1995}
Suppose that $\sigma>0$ and $t>0$. If $f\in H^1(\R)$, then
\begin{equation*}\label{eq4}
\lim_{t\to0^+}K_{\sigma}(f,t)_{H^1}=0.
\end{equation*}
\end{proposition}



\begin{definition}\label{def3}\cite[Page 176]{Lu1995}
Suppose that $m\in L^\infty(\R)$. If a family of operators $\{M_\epsilon\}_{\epsilon>0}$ defined by the equality
\begin{equation}\label{eq4}
\widehat{M_{\epsilon} f} (x)=m(\epsilon x)\widehat{f}(x),\quad f\in H^1 \cap L^2(\R)
\end{equation}
can be extended into a family of bounded operators on $H^1(\R)$, and their norms are uniformly bounded in $\epsilon$, the $m(x)$ is called a {\it homogeneous $H^1$ multiplier}.
\end{definition}

\section{Main results \label{s3}}

In this section, we first prove that the partial Hausdorff integrals $\{F_\epsilon\}_{\epsilon>0}$ are uniformly bounded on $H^1(\R)$, then we further obtain the approximation via partial Hausdorff integrals on $H^1(\R)$.

\begin{theorem}\label{th2}
Suppose that $a$ is a measurable function and $\varphi\in L^1(\R)$ satisfies $\int_{\R}\varphi(t)dt=1$ and
\begin{equation}\label{eq1103}
\varphi|a|^\frac{1}{2}\in L^1(\R).
\end{equation}
Then for any $\epsilon>0$ and $f\in H^1(\R)$,
\[
\|F_\epsilon\|_{H^1}\leq C\|f\|_{H^1},
\]
where $C$ is independent of $\epsilon$.
\end{theorem}

To prove Theorem \ref{th2}, we need two lemmas as follows.

\begin{lemma}\label{lem2}\cite[Page 114]{Stein1993}
Suppose \( g \) is a locally integrable function away from the origin on $\R$, and $|\widehat{g}(x)|\leq A_1$, $x\in\R$. Let
\begin{equation}\label{eq6}
(Tf)(x):=f*g(x), ~f\in L^2(\R).
\end{equation}
If 
\begin{equation}\label{eq7}
\int_{|x|\geq2|y|}|g(x-y)-g(x)|dy\leq A_2,~\text{whenever}~y\neq 0,
\end{equation}
then $T$ is bounded on $H^1(\R)$, that is,
\begin{equation}\label{eq8}
\|Tf\|_{H^1}\leq C(A_1,A_2) \|f\|_{H^1},~f\in H^1(\R).
\end{equation}
\end{lemma}

\begin{lemma}\label{lem3}\cite[Chapter 6]{Bracewell1965}
Let $f(x)=\frac{\sin x}{x},x\neq0$. The Fourier transform of $f$ is
\[
h(x):=\widehat{f}(x)
=
\begin{cases}
\pi, &|x|<1;\\
\frac{1}{2}\pi, &|x|=1;\\
0, &|x|>1.
\end{cases}
\]
\end{lemma}

\begin{proof}[Proof of Theorem \ref{th2}]
 Let $f\in L^2(\R)$. By substituting $\epsilon s=\tilde{s}$ and Fubini's theorem, we have
\[
F_\epsilon(x)=\frac{1}{\pi}\int_{\R} f(x-s) \int_{\R} \varphi(t) \frac{\sin(|a(t)|\epsilon^{-1}s)}{s}dtds.
\]
Denote
\[
K_\epsilon (s):=\frac{1}{\pi}\int_{\R} \varphi(t) \frac{\sin(|a(t)|\epsilon^{-1}s)}{s} dt,\quad K(s):=\frac{1}{\pi}\int_{\R} \varphi(t) \frac{\sin(|a(t)|s)}{s} dt.
\]
Then we get
\begin{equation}\label{eq16}
F_\epsilon (x)=K_\epsilon*f(x).
\end{equation}

Now let us prove that $K_\epsilon\in L^2(\R)$. By Minkowskis' inequality, substituting $|a(t)|\epsilon^{-1}s=\tilde{s}$, $\frac{\sin x}{x}$ is even function on $\R$, $|\sin x|\leq 1$, $|\sin x|\leq |x|$  and $(\ref{eq1103})$, we have
\begin{align}
\|K_\epsilon\|_{L^2}
&\leq
\frac{1}{\pi}\int_{\R}|\varphi(t)|\left(\int_{\R}\left|\frac{\sin(|a(t)|\epsilon^{-1}s)}{s}\right|^2ds\right)^\frac{1}{2}dt\nonumber\\
&=\frac{\epsilon^{-\frac{1}{2}}}{\pi}\int_{\R}|\varphi(t)||a(t)|^\frac{1}{2}dt\left(\int_{\R}\left|\frac{\sin s}{s}\right|^2ds\right)^\frac{1}{2}\nonumber\\
&= \frac{\epsilon^{-\frac{1}{2}}}{\pi}\int_{\R}|\varphi(t)||a(t)|^\frac{1}{2}dt\left(2\int_{0}^1\left|\frac{\sin s}{s}\right|^2ds+2\int_1^\infty\left|\frac{\sin s}{s}\right|^2ds\right)^\frac{1}{2}\nonumber\\
&\leq \frac{\epsilon^{-\frac{1}{2}}}{\pi}\int_{\R}|\varphi(t)||a(t)|^\frac{1}{2}dt\left(2\int_{0}^11ds+2\int_1^\infty\frac{1}{s^2}ds\right)^\frac{1}{2}\nonumber\\
&<\infty.
\end{align}
By the properties of the Fourier transform and substituting $\epsilon^{-1}y=\tilde{y}$ , we obtain
\begin{align}
\widehat{F_\epsilon}(x)
&=\widehat{K_\epsilon}(x)\widehat{f}(x)\nonumber\\
&=\widehat{f}(x) \int_{\R} \frac{1}{\pi} \int_{\R} \varphi(t) \frac{\sin(|a(t)|\epsilon^{-1}y)}{y} dt e^{-ixy}dy\nonumber\\
&=\widehat{f}(x) \int_{\R} \frac{1}{\pi} \int_{\R} \varphi(t) \frac{\sin(|a(t)|y)}{y} dt e^{-ix\epsilon y}dy\nonumber\label{eq12011}\\
&=\widehat{K}(\epsilon x)\widehat{f}(x).
\end{align}
By substituting $|a(t)|y=\tilde{y}$, Fubini's theorem and Lemma \ref{lem3}, we have
\begin{align}
\widehat{K}(x)
&=\int_{\R} \frac{1}{\pi} \int_{\R} \varphi(t) \frac{\sin(|a(t)|y)}{y} dt e^{-ixy}dy\nonumber\\
&=\frac{1}{\pi}  \int_{\R}\varphi(t)\int_{\R}\frac{\sin y}{y} e^{-ix|a(t)|^{-1}y}dydt \nonumber\\
&=\frac{1}{\pi} \int_{\R}\varphi(t)\left(\frac{\sin\cdot}{\cdot}\right)\widehat{~}(x|a(t)|^{-1})dt\nonumber 
\\
&=\frac{1}{\pi} \int_{\R} \varphi(t) h(x|a(t)|^{-1}) dt.\label{eq14} 
\end{align}
By $(\ref{eq14})$ and Lemma \ref{lem3}, we obtain
\begin{equation}\label{eq107}
\left|\widehat{K}(x)\right|\leq \|\varphi\|_{L^1(\R)}.
\end{equation}
The same steps as above can be used to obtain 
\begin{equation}\label{eq15}
\left|\widehat{K_\epsilon}(x)\right|\leq \|\varphi\|_{L^1(\R)}.
\end{equation}

For any $y\neq0$ and $|x|\geq 2|y|$, we have
\begin{equation}\label{eq13}
|x-y|\geq |x|-|y|\geq \frac{|x|}{2}.
\end{equation}
 Then by $(\ref{eq13})$, we get
\begin{align}
&\int_{|x|\geq 2|y|} \left|K_\epsilon(x-y)-K_\epsilon(x)\right|dy\nonumber\\
=&
\frac{1}{\pi}\int_{|x|\geq 2|y|} \left| \int_{\R} \varphi(t) \left[\frac{\sin(|a(t)|\epsilon^{-1}(x-y))}{x-y}-\frac{\sin(|a(t)|\epsilon^{-1}x)}{x}\right] dt \right| dy
\nonumber\\
\leq&
\frac{\|\varphi\|_{L^1(\R)}}{\pi}\int_{|x|\geq 2|y|}\frac{1}{|x-y|}dy+\frac{\|\varphi\|_{L^1(\R)}}{\pi}\int_{|x|\geq 2|y|}\frac{1}{|x|}dy
\nonumber\\
\leq&
\frac{\|\varphi\|_{L^1(\R)}}{\pi}\int_{|y|\leq\frac{|x|}{2}}\frac{2}{|x|}dy+\frac{\|\varphi\|_{L^1(\R)}}{\pi}\int_{|y|\leq\frac{|x|}{2}} \frac{1}{|x|}dy
\nonumber\\
=&
\frac{3}{2\pi}\|\varphi\|_{L^1(\R)}.\label{eq17}
\end{align}
By Lemma \ref{lem2} with $(\ref{eq15})$, $(\ref{eq16})$ and $(\ref{eq17})$ we obtain
\begin{equation*}
\|F_\epsilon\|_{H^1(\R)}\leq C \|f\|_{H^1(\R)},~f\in H^1(\R),
\end{equation*}
where $C$ is independent of $\epsilon$.
\end{proof}

\begin{theorem}\label{th1}
Let $\varphi$ and $a$ be as in Theorem \ref{th2}. Additionally, assume that $a$ is odd and such that $|a|$ is decreasing, positive and bijective on $(0,+\infty)$, and that for every set $E$ of zero Lebesgue measure, the set $a^{-1}(E)$ also has zero measure. Then for $\sigma>0$ and any $f\in H^1(\R)$, we have
\begin{equation}\label{eq10}
\|F_\epsilon -f\|_{H^1}\leq C K_{\sigma}(f,\epsilon)_{H^1}\to0~(\epsilon\to0^+),
\end{equation}
where $C>0$ is independent of $\epsilon$.
\end{theorem}

To prove Theorem \ref{th1}, we need a lemma as follows. 

\begin{lemma}\label{lem1}\cite[Page 179]{Lu1995}
Suppose that $\sigma>0$ and $m(x)$ is a homogeneous $H^1(\R)$ multiplier, and the family of operators $\{ M_\epsilon \}_{\epsilon>0}$ is defined by $(\ref{eq4})$. If there exists a $d>0$ such that
\begin{enumerate}
\item[(1)] $|m(x)-1|\leq C |x|^\sigma$ for $|x|\leq d$;
\item[(2)] For each $0<R<d$, we have
\begin{equation}\label{eq5}
\int_{R/2<|x|<R} \left| m'(x) \right|^2 dx \leq C R^{2\sigma-1},
\end{equation}
\end{enumerate}
then 
\begin{equation}\label{eq6}
\|M_\epsilon f-f\|_{H^1}\leq C K_\sigma (f,\epsilon)_{H^1},~ f\in H^1(\R),
\end{equation}
where $C>0$ is independent of $R$.
\end{lemma}


\begin{proof}[Proof of Theorem \ref{th1}]
By $(\ref{eq12011})$, $(\ref{eq107})$ and Theorem \ref{th2}, we obtain that $\widehat{K}$ is a homogeneous $H^1$ multiplier. By Lemma \ref{lem3} and $(\ref{eq14})$, we get
\begin{equation}\label{eq18}
\widehat{K}(0)=\int_{\R} \varphi(t) dt=1.
\end{equation}
By Lebesgue's dominated convergence theorem with $h'(x)=0$, a.e. $x\in\R$ and $a^{-1}(E)$ has zero Lebesgue measure for any zero measure set $E\subset\R$, we obtain
\begin{equation}\label{eq11281}
[\widehat{K}(x)]'=0.
\end{equation}
From this and Lagrange's mean value theorem, it follows that
\begin{equation}\left|\widehat{K}(x)-1\right|
=
\left|\widehat{K}(x)-\widehat{K}(0)\right|
=0
\leq
|x|^\sigma.\label{eq19}
\end{equation}
Therefore, by Lemma \ref{lem1} with $(\ref{eq19})$ and $(\ref{eq11281})$, we obtain
\[
\|F_\epsilon -f\|_{H^1}\leq C K_{\sigma}(f,\epsilon)_{H^1},~f\in H^1.
\]
Consequently, by Proposition \ref{pro1}, we have
\[
\|F_\epsilon -f\|_{H^1}\leq C K_{\sigma}(f,\epsilon)_{H^1}\to0~(\epsilon\to0^+), ~f\in H^1.
\]
\end{proof}

\section{Examples \label{s4}}
In the results concerning partial Hausdorff integrals, there is flexibility in the choice of \(\varphi\). This section will select some types of typical specific functions for \(\varphi\).\begin{example}\label{ex1}
Let $p>1$, $\sigma>0$, $\epsilon>0$ and $f\in L^2(\R)$. When $a(t)=\frac{1}{t}$ and $\varphi(t)=\frac{p-1}{2|t|^p}\chi_{(1,+\infty)}(|t|)$, the general Hausdorff operator $H_{\varphi,a}$ reduces to
\[
H_{\varphi,a}(f)(x)=\frac{p-1}{2}\int_{|t|>1}\frac{1}{|t|^{p+1}}f\left(\frac{x}{t}\right) dt,~x\in\R.
\]
The corresponding partial Hausdorff integrals reduces to
\[
F_\epsilon=\frac{p-1}{2\pi}\int_{|t|>1}\frac{1}{|t|^p}\int_{\R}f(x-\epsilon s)\frac{\sin\left(\frac{s}{|t|}\right)}{s}dsdt,~x\in\R.
\]
Obviously, $\varphi(t)$ and $a(t)$ satisfy the assumptions of Theorem \ref{th2} and Theorem \ref{th1}, thus for $f\in H^1(\R)$
\[
\|F_\epsilon\|\leq C_1\|f\|_{H^1}~\text{and}~\|F_\epsilon-f\|_{H^1}\leq C_2K_\sigma(f,\epsilon)\to0,~(\epsilon\to0^+),
\] 
where $C_1$, $C_2>0$ are independent of $\epsilon$.
\end{example}

\begin{example}\label{ex2}
Let $p<\frac{1}{2}$, $\sigma>0$, $\epsilon>0$ and $f\in L^2(\R)$. When $a(t)=\frac{1}{t}$ and $\varphi(t)=\frac{1-p}{2|t|^p}\chi_{(0,1)}(|t|)$, the general Hausdorff operator $H_{\varphi,a}$ reduces to
\[
H_{\varphi,a}(f)(x)=\frac{1-p}{2}\int_{|t|<1}\frac{1}{|t|^{p+1}}f\left(\frac{x}{t}\right)dt,~x\in\R.
\]
The corresponding partial Hausdorff integrals reduces to
\[
F_\epsilon=\frac{1-p}{2\pi}\int_{|t|<1}\frac{1}{|t|^p}\int_{\R}f(x-\epsilon s)\frac{\sin\left(\frac{s}{|t|}\right)}{s}dsdt,~x\in\R.
\]
Obviously, $\varphi(t)$ and $a(t)$ satisfy the assumptions of Theorem \ref{th2} and Theorem \ref{th1}, thus for $f\in H^1(\R)$
\[
\|F_\epsilon\|_{H^1}\leq C_1\|f\|_{H^1}~\text{and}~\|F_\epsilon-f\|_{H^1}\leq C_2K_\sigma(f,\epsilon)\to0,~(\epsilon\to0^+),
\] 
where $C_1$, $C_2>0$ are independent of $\epsilon$.
\end{example}

\begin{example}
In Example \ref{ex2}, if we pick \(p = 0\), the operator \(H_{\varphi,a}\) reduces to the adjoint Hardy operator \(H^*\),
\[
H^*(f)(x):=\frac{1}{2}\int_{|t|>|x|}\frac{f(t)}{|t|}dt,~x\in\R.
\]
\end{example}

\begin{example}\label{ex3}
Let $\alpha>0$, $\sigma>0$, $\epsilon>0$ and $f\in L^2(\R)$. When $a(t)=\frac{1}{t}$ and $\varphi(t)=\frac{1}{2}(1+\alpha)(1-|t|)^\alpha\chi_{(0,1)}(|t|)$, the general Hausdorff operator $H_{\varphi,a}$ reduces to Riemann-Liouville type integral, i.e.
\[
H_{\varphi,a}(f)(x)=\frac{1+\alpha}{2}\int_{|t|<1}\frac{(1-|t|)^\alpha}{|t|}f\left(\frac{x}{t}\right)dt,~x\in\R.
\]
The corresponding partial Hausdorff integrals reduces to
\[
F_\epsilon=\frac{1+\alpha}{2\pi}\int_{|t|<1}(1-|t|)^\alpha \int_{\R}f(x-\epsilon s)\frac{\sin\left(\frac{s}{|t|}\right)}{s}dsdt,~x\in\R.
\]
Obviously, $\varphi(t)$ and $a(t)$ satisfy the assumptions of Theorem \ref{th2} and Theorem \ref{th1}, thus for $f\in H^1(\R)$
\[
\|F_\epsilon\|_{H^1}\leq C_1\|f\|_{H^1}~\text{and}~\|F_\epsilon-f\|_{H^1}\leq C_2K_\sigma(f,\epsilon)\to0,~(\epsilon\to0^+),
\] 
where $C_1$, $C_2>0$ are independent of $\epsilon$.
\end{example}

\noindent{\bf Funding } This project is supported by National Natural Science Foundation of China (No.12261083) and Natural Science Foundation of Xinjiang Province (No.2024D01C40).

\noindent{\bf Data Availability} No datasets were generated or analysed during the current study.

\section*{Declarations}

\noindent{\bf Ethical Approval} Not applicable.

\noindent{\bf Competing interests} The authors declare no competing interests.

\end{document}